A cannon is shot off a cliff at 35 degree angle off a cliff of 4 m in height. What is the x of the location where the cannon ball land if its velocity is 7.03m/s
Answer #1

VI

vi=7.03m.s

vix=vicosθ=7.03cos35 = 5.76 m/s

viy=visinθ=7.03sin35 = 4.03 m/s

Let us first find the time of flight of the cannon,

Use kinematic eqn,

y= viy*t -0.5gt^2

4.0 = 4.03*t -0.5*9.8*t^2

Calculations give t=1.4s

Range x = vix*t = 5.76*1.4 =

x = 8.06 m